OpenStudy (anonymous):

how do i solve log(3x+5)=2?

OpenStudy (anonymous):

do you know how to get rid of a log?

OpenStudy (anonymous):

like make it ln?

OpenStudy (anonymous):

if you raise 10 to the log of some number, it just equals that number

OpenStudy (anonymous):

ohhhh well i dont have a calculator with me...

OpenStudy (anonymous):

you don't need one

OpenStudy (anonymous):

okay well i have no idea how to solve it..

OpenStudy (anonymous):

10^(logx) = x

OpenStudy (anonymous):

10^(log(x+2)) = x+2

OpenStudy (anonymous):

10^2 = 100

OpenStudy (anonymous):

so how would i do log(3x+5)=2?

OpenStudy (anonymous):

10^(log(3x+5)) = 10^2

OpenStudy (anonymous):

what does ^ mean? lol

OpenStudy (anonymous):

raise to a power

OpenStudy (anonymous):

so 10^2 means 10 squared

OpenStudy (anonymous):

ohhhh!

OpenStudy (anonymous):

okay so then what would the answer be? cause idk how to evaluate it without a calc.

OpenStudy (anonymous):

3x+5 = 100

OpenStudy (anonymous):

ohhh okay thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so would it be the same thing for log160-log5?

OpenStudy (anonymous):

it wouldn't be, but look wher you posted the question for the answer